ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 193, 2013

Aizenbud, A. A partial analog of the integrability theorem for distributions on p-adic spaces and applications

233

Bartholdi, L. Branch rings, thinned rings, tree enveloping rings — Erratum

507

Bayart, F. and G. Costakis Cyclic operators with finite support  131

Bourgain, J. Moment inequalities for trigonometric polynomials with spectrum in curved hypersurfaces

441

Brešar, M. and I. Klep A local-global principle for linear dependence of noncommutative polynomials

71

Brooks, S. and E. Lindenstrauss Non-localization of eigenfunctions on large regular graphs

1

Brownlowe, N., A. Sims and S. T. Vittadello Co-universal C^{∗}- algebras associated to generalised graphs

399

Burger, M., N. Ozawa and A. Thom On Ulam stability  109

Costakis, G. See F. Bayart Cuenca Mira, J. A. Some classes of pre-Hilbert algebras with norm-

one central idempotent

343

Delzant, T. and L. Potyagailo On the complexity and volume of hyperbolic 3-manifolds

209

Ferreira, L. C. F. and M. Montenegro A Fourier approach for nonlinear equations with singular data

83

Foreman, M. Calculating quotient algebras of generic embeddings  309

Giordano, P. and M. Kunzinger Topological and algebraic structures on the ring of Fermat reals

459

Goodrick, J., B. Kim and A. Kolesnikov Amalgamation functors and boundary properties in simple theories

169

Grynkiewicz, D. J., A. Philipp and V. Ponomarenko Arithmetic-progression-weighted subsequence sums

359

Jouve, F., E. Kowalski and D. Zywina Splitting fields of characteristic polynomials of random elements in arithmetic groups

263

Kim, B. See J. Goodrick Klep, I. See M. Brešar Kolesnikov, A. See J. Goodrick Kowalski, E. See F. Jouve Kunzinger, M. See P. Giordano Lindenstrauss, E. See S. Brooks

Minsky, Y. N. On dynamics of Out(F_{n}) on PSL_{2}(ℂ) characters  47

Montenegro, M. See L. C. F. Ferreira Ozawa, N. See M. Burger Philipp, A. See D. J. Grynkiewicz Ponomarenko, V. See D. J. Grynkiewicz Potyagailo, L. See T. Delzant Schmidt, T. On unitary deformations of smooth modular representa-

tions

15

Sims, A. See N. Brownlowe Thom, A. See M. Burger Vittadello, S. T. See N. Brownlowe Zywina, D. See F. Jouve